D e f i n e an e q u i l a t e r a l q u a d r i l a t e r a l.

D e f i n e  an e q u i l a t e r a l q u a d r i l a t e r a l.

Grade:11

4 Answers

Rituraj Tiwari
3 years ago
An equilateral quadrilateral, i.e. the one with all sides equal, is a rhombus. In a square, rectangle, or rhombus, the opposite side lines are parallel. A quadrilateral with the opposite side lines parallel is known as a parallelogram.
Janarth
3 years ago
hi!  An equilateral quadrilateral, i.e. the one with all sides equal, is a rhombus. In a square, rectangle, or rhombus, the opposite side lines are parallel. A quadrilateral with the opposite side lines parallel is known as a parallelogram.thanks.
Ram Kushwah
3 years ago
Equilateral qudrilateral:
We can call it a regular quadrilateral
for a regular quadrilateral n=4
thus each enagle=(2*n-4)*90/n
=(2*4-4)*90/4
=4*90/4
= 90°
So each angle of the quadrilateral is 90°
and all the sides are equal 
Such quadrilateral ia called a square 
Thus square is only an equilateral 
 
Mahatma Patel
2 years ago
An equilateral quadrilateral i.e. the one with all sides equal,is a rhombus. In a square, rectangle or a rhombus, the opposite side lines are parallel. A quadrilateral with opposite side lines are parallel is known as parallelogram.